DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, December 6

Thompson 331 2:30 pm

NOTE: Special day, time and room

On the concentration of blow-up solutions of some nonlinear

dispersive equations

Luis Vega

(Universidad del Pais Vasco, Spain)